# Quantifying Memorization of Domain-Specific Pre-trained Language Models using Japanese Newspaper and Paywalls

### v1

@@ Abstract
Dominant pre-trained language models (PLMs) have been successful in high-quality natural language generation. However, the analysis of their generation is not mature: do they acquire generalizable linguistic abstractions, or do they simply memorize and recover substrings of the training data? Especially, few studies focus on domain-specific PLM. In this study, we pre-trained domain-specific GPT-2 models using a limited corpus of Japanese newspaper articles and quantified memorization of training data by comparing them with general Japanese GPT-2 models. Our experiments revealed that domain-specific PLMs sometimes “copy and paste” on a large scale. Furthermore, we replicated the empirical finding that memorization is related to duplication, model size, and prompt length, in Japanese the same as in previous English studies. Our evaluations are relieved from data contamination concerns by focusing on newspaper paywalls, which prevent their use as training data. We hope that our paper encourages a sound discussion such as the security and copyright of PLMs.
Quantifying Memorization of Domain-Specific Pre-trained Language Models using Japanese Newspaper and Paywalls
Shotaro Ishihara Nikkei Inc. Tokyo, Japan shotaro.ishihara@nex.nikkei.com
@@ 1 Introduction
Pre-trained language models (PLMs) have shown great capabilities in solving various tasks in natural language processing (Yang et al., 2023; Zhao et al., 2023). Statistical language models learn the probability of word occurrence, and pre-training on large datasets for large neural networks has become popular. This extension has led to fluent natural language generation and has been reported to perform well when fine-tuned for many downstream tasks (Radford et al., 2018). For much larger models called large language models (LLMs), downstream tasks can be solved without parameter updates (Radford et al., 2019; Brown et al., 2020). Social recognition such as ChatGPT [^1] https://openai.com/blog/chatgpt is steadily increasing.
@@
As practical applications evolve, critical views on the generation of PLMs are becoming apparent in security and copyright (Bender et al., 2021; Bommasani et al., 2021; Weidinger et al., 2022). Prior research has indicated that neural networks have the property of unintentionally memorizing and outputting the training data (Carlini et al., 2019, 2021, 2023; Lee et al., 2023). In particular, Carlini et al. (2021) demonstrated that memorized personal information (names, phone numbers, and email addresses) can be extracted from GPT-2 models (Radford et al., 2019). This can lead to an invasion of privacy, reduced utility, and reduced ethical practices (Carlini et al., 2023). If there is no novelty in the generation, there would be a problem in terms of copyright (McCoy et al., 2023; Franceschelli and Musolesi, 2023).
Despite its significance, this discussion remains in its infancy (Ishihara, 2023). Initial studies remain on the qualitative side (Carlini et al., 2021), and several studies have begun to focus on quantitative evaluations (Lee et al., 2022; Kandpal et al., 2022; Ippolito et al., 2022; Tirumala et al., 2022; Downey et al., 2022; Carlini et al., 2023; Lee et al., 2023). These studies were conducted mainly in English, and their reproducibility was uncertain under domain-specific conditions. Memorization of machine learning models is generally associated with overfitting (Yeom et al., 2018; Zhang et al., 2021a), which is even more important to discuss under conditions when it is difficult to prepare large training data. Compared to general corpora, considerations of security and copyright are increasingly important for rare corpora.
This study is the first attempt to quantify the memorization of domain-specific PLMs using a limited corpus of Japanese financial newspaper articles. Our research objective is to identify trends in memorization of domain-specific PLMs. We argue that newspaper articles are suitable for evaluating the memorization of PLMs because their paywall characteristics prevent their use as training data (Section 2). First, we defined memorization and developed a framework for quantifying the memorization of domain-specific PLMs using Japanese newspaper articles (Section 3). Secondly, we pre-trained domain-specific GPT-2 models and observed that they sometimes memorized and output the training data on a large scale (Section 4). Experiments reported that memorization is related to duplication, model size, and prompt length. These empirical findings, which had been reported in previous studies in English, were found for the first time in Japanese. Finally, we discuss future research directions (Section 5).
@@ 2.1 Memorization of PLMs
Memorization of PLMs refers to the phenomenon of outputting fragments of the training data. Research on memorization is diverse, with various definitions and assumptions. This study focuses on autoregressive language models, such as the GPT family Radford et al. (2018, 2019); Brown et al. (2020); Black et al. (2022). These are promising models as of 2023.
@@
Our study designed the first of these definitions in Japanese and reported the experimental results. Both definitions of memorization are ambiguous in languages without obvious token delimiters such as Japanese. Definitions based on the concepts of differential privacy (Jagielski et al., 2020; Nasr et al., 2021) and counterfactual memorization (Zhang et al., 2021b) are beyond the scope of this study.
2.2 Issues with Memorization of PLMs
There are discussions of issues such as security and copyright with the memorization of PLM. Our study of quantifying memorization serves to confront these issues precisely.
Training data extraction.
Training data extraction is a security attack related to the memorization of PLMs (Ishihara, 2023). Many studies follow the pioneering work of Carlini et al. (2021). They reported that a large amount of information could be extracted by providing GPT-2 models with a wide variety of prompts (generating candidates) and performing membership inference (Shokri et al., 2017). In particular, when dealing with PLMs with sensitive domain-specific information such as clinical data, the leakage of training data can lead to major problems (Nakamura et al., 2020; Lehman et al., 2021; Jagannatha et al., 2021; Singhal et al., 2022; Yang et al., 2022). It is necessary to discuss from the perspective of human rights, such as the right to be forgotten (Li et al., 2018; Ginart et al., 2019; Garg et al., 2020), in terms of the unintentional accumulation and extraction of personal information (Henderson et al., 2022).
Novelty in text generation.
There has been a traditional research area for evaluating the quality of text generation, but few studies have focused on novelty. McCoy et al. (2023) emphasize that the research community should focus on novelty as well as fluency (Mutton et al., 2007), factual accuracy (Kryscinski et al., 2020), and diversity (Zhu et al., 2018; Hashimoto et al., 2019). Novelty in text generation is directly related to the discussion of copyright (Franceschelli and Musolesi, 2023). Lee et al. (2023) analyzed plagiarism patterns in PLMs using English domain-specific corpora.
2.3 Quantifying Memorization of PLMs
Recent studies have quantitatively evaluated memorization related to these issues (Lee et al., 2022; Kandpal et al., 2022; Ippolito et al., 2022; Tirumala et al., 2022; Downey et al., 2022; Carlini et al., 2023; Lee et al., 2023; McCoy et al., 2023).
@@ 2.3 Quantifying Memorization of PLMs
According to the first comprehensive quantitative studies (Carlini et al., 2023), the memorization of PLMs is strongly related to the training set string duplications, model size, and prompt length. In particular, there are related reports on the association of string duplications in the training set with the memorization of PLMs (Tirumala et al., 2022; Lee et al., 2023). Carlini et al. (2023) used the variation of the definition of eidetic memorization, and Ippolito et al. (2022) confirmed similar results with the definition of approximate memorization.
Our study examines domain-specific PLMs using Japanese financial newspaper articles. This is the first domain-specific study of PLMs in a non-English language, although there are some English examples. If the data size is domain-specific and small, people tend to train PLMs with multiple epochs. However, increasing the number of epochs is equivalent to string duplications, and its effect on memorization should be particularly considered.
Construction of evaluation sets.
@@
We describe the quantification methods used in the pioneering study (Carlini et al., 2023) and point out the potential for improvement. Owing to inference time limitations, it is not possible to evaluate memorization using all of the training data. For example, Carlini et al. (2023) targeted GPT-Neo models (Black et al., 2022) and constructed an evaluation set by sampling 50,000 samples from the Pile dataset (Gao et al., 2020) used for pre-training. Sampling and string splitting are unavoidable during the construction of the evaluation set, as shown in Figure 1. They assumed that the distribution of string duplicates was related to the memorization of PLMs. Each sampled sentence was divided into prompts of each length from 50 to 500 tokens at the beginning, with the following 50 tokens as references.
However, this splitting does not consider the importance of references. In other words, it does not consider whether references are protected subjects against security concerns.
2.4 Newspaper Paywalls as Evaluation Sets
@@
We argue that the use of newspaper articles can benefit the construction of evaluation sets. Newspaper paywall refers to a method of restricting access to online content through a paid subscription (Myllylahti, 2016). Online news services with paid subscription plans often publish newspaper articles only at the beginning, with the rest of the text available only to their members. This system creates a real-world setting in which there is a private part following the public part as illustrated in Figure 2. The use of private parts as references can achieve the splitting in which publishers hide important information that they want to preserve.
Figure 1: The existing method for constructing an evaluation set for quantifying memorization. This procedure requires sampling data from the training set used to pre-train and splitting the text into prompts and references.
We also present that newspaper paywalls can provide a solution to data contamination. The memorization of PLMs has been identified as damaging the integrity of the evaluation set. Several studies have identified the inclusion of evaluation sets in the large datasets used for pre-training, which has led to unfairly high performance (Magar and Schwartz, 2022; Jacovi et al., 2023; Aiyappa et al., 2023). In contrast, some parts of newspaper articles are available only to paying subscribers. This ensures that they are not used for training PLMs with common web datasets. This has significant value for the accurate evaluation of the memorization of PLMs.
@@
Newspaper paywalls are often discussed in the literature tied to journalism. For example, Kim et al. (2020) examined the impact of newspaper paywalls on daily page views and differences among publishers. Several other studies were conducted in the context of publishers’ digital strategies (Myllylahti, 2014; Carson, 2015; Sjøvaag, 2016). This study assigns new roles to newspaper paywalls. Newspaper articles are widespread in many languages; therefore, our proposal has the appeal of high versatility in low-resource languages.
3 Methodology for Quantification
@@
Figure 2: The procedure of quantifying the memorization of PLMs in this study. First, we pre-trained GPT-2 models using newspaper articles as a training set. We then generated strings using the public part as a prompt. Finally, the memorization was quantified using the private part.
Figure 4: Histogram of the number of characters up to the end of the first sentence in the private part in the evaluation set. Nine articles exceeded 200 characters and were therefore skipped in the visualization.
This section explains the problems addressed in this study. Specifically, we first design definitions of memorization in Japanese and then construct an evaluation set using newspaper paywalls. Finally, we describe the procedure for quantifying the memorization of PLMs (Figure 2), following a previous study (Carlini et al., 2021).
3.1 Definitions of memorization in Japanese
@@
This study designed two definitions of memorization, as described in Section 2.1. While previous studies were based on English words, we must consider that there are no spaces between words in Japanese. The definitions of the Japanese memorization of PLMs in this study are designed as follows.
eidetic memorization
@@
is measured by the number of forward-matching characters. This is a definition that is independent of the properties of the word segmenter and tokenizer. Therefore, it has advantages in dealing with languages without explicit word boundaries, such as Japanese. As this study uses Japanese newspaper articles and its paywall, we had to use a derivation that is indeed slightly different from the original eidetic memorization. It is a derivation of the original definition with the restriction of forward-matching characters.
approximate memorization
@@
is measured by a normalized Levenshtein distance (Yujian and Bo, 2007). The Levenshtein distance is a measure of the number of characters required to match one string to the other. We convert this value to similarity by dividing it by the number of characters of the higher value.
3.2 Construction of Evaluation Sets
As a dataset containing information on newspaper paywalls, we selected the corpus of Japanese financial newspaper articles provided by Nikkei Inc [^2] https://aws.amazon.com/marketplace/seller-profile?id=c8d5bf8a-8f54-4b64-af39-dbc4aca94384. The newspaper articles were covered from March 23, 2010333Launch date of Nikkei’s online edition to December 31, 2021. Note, that this corpus was filtered to include approximately one billion (B) tokens. In this corpus, the shorter of the first 200 words or half the number of words in the entire article is defined as the public part. Note that there are cases in which the entire article, including the private part, is made public according to various circumstances such as the importance of the topics.
We randomly sampled 1000 articles published in 2021 as our evaluation set. A histogram of the number of characters in the public part in the constructed evaluation set is shown in Figure 3. Most articles were approximately 200 words; however, some were shorter. Only a minority (25 articles) ended the public part using punctuation marks [^4] Japanese punctuation mark is “。”.. The private parts are extremely long for some articles, and we extracted them until the end of the first sentence [^5] We used bunkai (https://github.com/megagonlabs/bunkai). to simplify the problem (Figure 4).
3.3 Procedure of Quantification
In this study, we attempted to quantify the memorization of PLMs using a procedure similar to that in Carlini et al. (2023) (Figure 2). First, for preparation, GPT-2 models were pre-trained on all sentences in both the public and private parts of the articles. For a given article in the evaluation set, we considered the string in the public part to be a prompt and generated a string that follows. We generated a single string from a single prompt using a greedy method that produced the word with the highest conditional probability each time. The choice of decoding strategy is a matter for future studies as described in Section 5. Finally, the degree of memorization is evaluated by comparing the generated string with the private part. We used the two Japanese definitions of memorization defined in Section 3.1.
@@ 3.3 Procedure of Quantification
4 Experiments
This section reports our findings from experiments under various conditions. First, multiple PLMs were prepared, and then memorization was quantified. We analyzed the results from a quantitative and qualitative perspective.
4.1 Preparation of PLMs
We used both domain-specific and general GPT-2 models in our experiments for comparison.

### v2

# Quantifying Memorization and Detecting Training Data of Pre-trained Language Models using Japanese Newspaper

@@ Abstract
Dominant pre-trained language models (PLMs) have demonstrated the potential risk of memorizing and outputting the training data. While this concern has been discussed mainly in English, it is also practically important to focus on domain-specific PLMs. In this study, we pre-trained domain-specific GPT-2 models using a limited corpus of Japanese newspaper articles and evaluated their behavior. Experiments replicated the empirical finding that memorization of PLMs is related to the duplication in the training data, model size, and prompt length, in Japanese the same as in previous English studies. Furthermore, we attempted membership inference attacks, demonstrating that the training data can be detected even in Japanese, which is the same trend as in English. The study warns that domain-specific PLMs, sometimes trained with valuable private data, can “copy and paste” on a large scale. [^1] An early version of this study was accepted for non-archival track of the Fourth Workshop on Trustworthy Natural Language Processing (Ishihara, 2024).
Quantifying Memorization and Detecting Training Data of Pre-trained Language Models using Japanese Newspaper
Shotaro Ishihara Nikkei Inc. Tokyo, Japan shotaro.ishihara@nex.nikkei.com                        Hiromu Takahashi Independent Researcher Tokyo, Japan hiromu.takahashi56@gmail.com
@@ 1 Introduction
As pre-trained language models (PLMs) have become increasingly practical, critical views on the memorization of PLMs are emerging in security and copyright (Bender et al., 2021; Bommasani et al., 2021; Weidinger et al., 2022). Prior research has indicated that neural networks have the property of unintentionally memorizing and outputting the training data (Carlini et al., 2019, 2021, 2023; Lee et al., 2023; Yu et al., 2023). In particular, Carlini et al. (2021) demonstrated that memorized personal information can be detected from GPT-2 models (Radford et al., 2019). This can lead to an invasion of privacy, reduced utility, and reduced ethical practices (Carlini et al., 2023). If there is no novelty in the generation, there would be a problem with copyright (McCoy et al., 2023; Franceschelli and Musolesi, 2023).
Research on memorization of PLMs has been intensively advanced, and empirical findings have been reported (Ishihara, 2023). Initial studies remain on the qualitative side (Carlini et al., 2021), and subsequent studies have begun to focus on quantitative evaluations. According to one of the first comprehensive quantitative studies (Carlini et al., 2023), the memorization of PLMs is strongly related to the string duplications in the training set, model size, and prompt length. Benchmarking of memorized string detection has also progressed, including constructing evaluation sets Shi et al. (2024); Duarte et al. (2024); Kaneko et al. (2024); Duan et al. (2024).
These studies were conducted in English, and their reproducibility is uncertain under domain-specific conditions. Domain-specific PLMs are sometimes built on rare private corpora and have smaller pre-training corpora than general PLMs. When the data size is small, models tend to be pre-trained in multiple epochs. However, increasing the number of epochs is equivalent to string duplications, which risks increased memorization. Furthermore, security and copyright considerations become increasingly important, as the memorized contents tend to be more specific than general corpora. We, therefore, pose the following practically significant questions about domain-specific PLMs: how much of the pre-training data is memorized, and is the memorized data detectable?
This study is the first attempt to quantify the memorization of domain-specific PLMs using a limited corpus of Japanese newspaper articles. Our research objective is to identify the memorization properties of domain-specific PLMs. First, we developed a framework for quantifying the memorization and detecting training data of PLMs using Japanese newspaper articles (Section 3). We then pre-trained domain-specific GPT-2 models and quantified their memorization (Section 4). Furthermore, we addressed membership inference attacks (Shokri et al., 2017), which predicts whether the output string was included in the training data (Section 5).
The main findings and contributions of this paper are summarized as follows.
  • Quantification: Japanese PLMs were demonstrated to sometimes memorize and output the training data on a large scale. Experiments reported that memorization was related to duplication, model size, and prompt length. These empirical findings, which had been reported in English, were found for the first time in Japanese.
  • Detection: Experiments demonstrated that the training data was detected from PLMs even in Japanese. The membership inference approach suggested in English was successful with the AUC (area under the ROC curve) score of approximately 0.6. As well as the empirical findings of memorization, the more duplicates and the longer the prompt, the easier the detection was.
@@ 2.1 Memorization of PLMs
Memorization of PLMs refers to the phenomenon of outputting fragments of the training data. Research on memorization is diverse, with various definitions and assumptions. We focus on autoregressive language models, such as the GPT family Radford et al. (2018, 2019); Brown et al. (2020); Black et al. (2022). These are promising models and major research targets.
@@
Our study designed the first of these definitions in Japanese and reported the experimental results. Both definitions of memorization are ambiguous in languages without obvious token delimiters such as Japanese. Definitions based on the concepts of differential privacy (Jagielski et al., 2020; Nasr et al., 2021) and counterfactual memorization (Zhang et al., 2023) are beyond the scope of this study.
Issues with memorization of PLMs.
Training data extraction is a security attack related to the memorization of PLMs (Ishihara, 2023). Many studies follow the pioneering work of Carlini et al. (2021). They reported that a large amount of information could be extracted by providing GPT-2 models with various prompts (generating candidates) and performing membership inference. In particular, when dealing with PLMs with sensitive domain-specific information such as clinical data, the leakage of training data can lead to major problems (Nakamura et al., 2020; Lehman et al., 2021; Jagannatha et al., 2021; Singhal et al., 2023; Yang et al., 2022). It is also necessary to discuss from the perspective of human rights, such as the right to be forgotten (Li et al., 2018; Ginart et al., 2019; Garg et al., 2020).
There has been a traditional research area for evaluating the quality of text generation, but few studies have focused on novelty (McCoy et al., 2023). Novelty in text generation is directly related to the discussion of copyright (Franceschelli and Musolesi, 2023). Lee et al. (2023) analyzed plagiarism patterns in PLMs using English domain-specific corpora.
The memorization of PLMs has also been identified as data contamination damaging the integrity of the evaluation set. Several studies have identified the inclusion of evaluation sets in the large datasets used for pre-training, which has led to unfairly high performance (Magar and Schwartz, 2022; Jacovi et al., 2023; Aiyappa et al., 2023).
Our study of quantifying memorization and performing membership inference would serve to confront these issues precisely in Japanese.
2.2 Quantifying Memorization and Detecting Training Data of PLMs
Recent studies have quantitatively evaluated memorization and related issues.
@@ 2.2 Quantifying Memorization and Detecting Training Data of PLMs
As mentioned in Section 1, empirical findings in English are known that the memorization of PLMs is strongly related to the string duplications in the training set, model size, and prompt length (Carlini et al., 2021). There are supportive reports for this finding for duplication (Lee et al., 2022; Tirumala et al., 2022; Lee et al., 2023; Ippolito et al., 2023; Kandpal et al., 2022; McCoy et al., 2023), model size Huang et al. (2022); Kandpal et al. (2022); Lee et al. (2023); Karamolegkou et al. (2023); Ippolito et al. (2023); McCoy et al. (2023), and prompt length Huang et al. (2022); Kandpal et al. (2022).
Evaluation sets for quantification.
@@
We describe the quantification methods used in the pioneering study (Carlini et al., 2023) and point out the potential for improvement. Owing to inference time limitations, it is impossible to evaluate memorization using all of the training data. For example, Carlini et al. (2023) targeted GPT-Neo models (Black et al., 2022) and constructed an evaluation set by sampling 50,000 samples from the Pile dataset (Gao et al., 2020) used for pre-training. Sampling and string splitting are unavoidable during the construction of the evaluation set, as shown in Figure 1. Each sampled sentence was divided into prompts of each length from 50 to 500 tokens at the beginning, with the following 50 tokens as references.
@@
However, this splitting does not consider the importance of references. In other words, it does not consider whether references are protected subjects against security concerns. We argue that using newspaper articles can provide real-world settings in data splitting via their paywalls. Newspaper paywall restricts access to online content through a paid subscription (Myllylahti, 2016). Online news services with paid subscription plans often publish newspaper articles only at the beginning, with the rest of the text available only to their members. This system creates a real-world setting in which there is a private part following the public part as illustrated in Figure 2. Using private parts as references can achieve the splitting in which publishers hide important information that they want to preserve.
@@
Newspaper paywalls are often discussed in the literature tied to journalism. For example, Kim et al. (2020) examined the impact of newspaper paywalls on daily page views and differences among publishers. Several other studies were conducted in the context of publishers’ digital strategies (Myllylahti, 2014; Carson, 2015; Sjøvaag, 2016).
Figure 1: The existing method for constructing an evaluation set for quantifying memorization and detecting training data. This procedure requires sampling data from the training set used to pre-train and splitting the text into prompts and references. Positive examples are created from training data and negative examples from new text that are guaranteed not to be training data.
Evaluation sets for training data detection.
To evaluate the detection of memorized training data from PLMs, it is necessary to have data that is guaranteed not to have been used for pre-training. A promising approach is to use new texts generated after constructing PLMs. Shi et al. (2024) constructed a dataset based on the creation date of the Wikipedia articles. Duarte et al. (2024) developed a dataset from the publication years of 165 books.
Along with evaluation sets, detection methods have been explored. For example, Shi et al. (2024) proposed Min-$k\%$ Prob, which extracts $k$ tokens with high log-likelihood and uses the average log-likelihood for detection. Min-$k\%$ Prob is regarded as one of the current prevailing methods (Kaneko et al., 2024; Zhang et al., 2024; Meeus et al., 2024). Kaneko et al. (2024) introduced SaMIA, which generates multiple candidates and calculates the average of the ROUGE-1 (Lin, 2004) without using the output of likelihood. The AUC score and TPR@10%FPR (True Positive Rate when False Positive Rate is 10 %) are used as the metrics (Mattern et al., 2023; Shi et al., 2024; Kaneko et al., 2024). Note that Carlini et al. (2022) recommended reporting TPR when FPR is low in membership inference assessments.
We use Japanese newspaper articles to construct the evaluation set and perform the existing detection method. Newspaper articles are generated daily, ensuring data is not used for pre-training. Given the widespread use of newspaper articles in many languages, our proposal has the appeal of high versatility in low-resource languages.
@@
Figure 2: The procedure of quantifying the memorization and training data detection of PLMs in this study. First, we pre-trained GPT-2 models using newspaper articles as a training set. We then generated strings using the public part as a prompt. The memorization was quantified using the private part. We also tackle the training data detection task, using articles used for pre-training as positive examples and not as negative examples.
3 Problem Statement & Methodology
This section explains the problem addressed in this study and the methodology (Figure 2). We use a methodology similar to that in Carlini et al. (2023).
3.1 Constructing Evaluation sets.
As described in Section 2, we construct evaluation sets using newspaper articles and paywalls.
Evaluation sets for quantification.
To quantify memorization, sentences need to be split into prompts and references. We propose to use the beginning of the newspaper article (the public part) as a prompt and the continuation in the paywall (the private part) as a reference.
Evaluation sets for training data detection.
Positive and negative examples are required to measure the performance of training data detection. We propose to use the newspaper articles used to construct the PLMs as positive examples and those published later as negative examples.
3.2 Quantifying Memorization
The three steps to quantify memorization are described.
Step 1. Preparing PLMs.
First, as a preparation, PLMs are built using all sentences containing both public and private parts of newspaper articles.
Step 2. Generating candidate.
For a given article in the evaluation set, we consider the string in the public part to be prompt and generate a string that follows.
Step 3. Calculating similarity.
@@ 3.2 Quantifying Memorization
The degree of memorization is evaluated by comparing the generated string with the private part. We designed two Japanese definitions of memorization of PLMs. While previous studies were based on English words, we must consider that there are no spaces between words in Japanese. The definitions of memorization in this study are as follows.
@@
  • The eidetic memorization is measured by the number of forward-matching characters. This is a definition that is independent of the properties of the word segmenter and tokenizer. Therefore, it has advantages in dealing with languages without explicit word boundaries, such as Japanese. As this study uses Japanese newspaper articles and their paywall, we had to use a derivation slightly different from the original eidetic memorization. It is a derivation of the original definition with the restriction of forward-matching characters.
@@
  • The approximate memorization is measured by a normalized Levenshtein distance (Yujian and Bo, 2007). The Levenshtein distance is a measure of the number of characters required to match one string to the other. We convert this value to similarity by dividing it by the number of characters of the higher value.
3.3 Detecting Training Data.
We also attempt to detect memorized training data. In this problem setting, there are two differences from quantifying memorization.
  • The reference is not available. This is because the situation where an attacker knows the reference is not realistic.
  • The likelihood of PLMs is available. We can get not only the output string but also the likelihood.
Therefore, instead of Step 3 in which memorization is quantified in terms of string similarity between the candidate and the reference, we establish Step 3’ in which membership probability is calculated.
Step 3’. Calculating membership probability.
For the detection method, we use Min-$k\%$ Prob for $k$ in $\{10,20,30,40,50,60\}$. As described in Section 2, Min-$k\%$ Prob calculates the membership probability by extracting and averaging $k$ tokens with high log-likelihood. The AUC score and TPR@10%FPR are reported in common with the previous studies.
@@
4 Experiment 1: Quantification
This section reports our findings from experiments under various conditions. First, multiple PLMs and the evaluation set were prepared, and then memorization was quantified. We analyzed the results from a quantitative and qualitative perspective.
4.1 Preparing Evaluation Set
As a dataset containing information on newspaper paywalls, we selected the corpus of Japanese newspaper articles provided by Nikkei Inc [^2] https://aws.amazon.com/marketplace/seller-profile?id=c8d5bf8a-8f54-4b64-af39-dbc4aca94384. The newspaper articles were covered from March 23, 2010333Launch date of Nikkei’s online edition to December 31, 2021. In this corpus, the shorter of the first 200 words or half the number of words in the entire article is defined as the public part. This corpus was filtered to include approximately 1-2 billion (B) tokens. Note that there are cases in which the entire article, including the private part, is made public according to various circumstances such as the importance of the topics.
We randomly sampled 1,000 articles published in 2021 as our evaluation set. The number of characters in the public part was approximately 200 words in most articles; however, some were shorter. Only a minority (25 articles) ended the public part using punctuation marks [^4] Japanese punctuation mark is “。”.. The private parts are extremely long for some articles, and we extracted them until the end of the first sentence [^5] We used bunkai (https://github.com/megagonlabs/bunkai). to simplify the problem. Histograms of the number of characters in the public and private part in the constructed evaluation set are shown in Figure 3 and 4.
Figure 4: Histogram of the number of characters up to the end of the first sentence in the private part of the evaluation set. Nine articles exceeded 200 characters and were therefore skipped in the visualization.
4.2 Step 1: Preparing PLMs
For comparison, we used both domain-specific and general GPT-2 models in our experiments.